A small box has a perimeter of 38 and a large box has a perimeter 0f 57. the width of the smaller box is 6 what is the width of the larger box?
Answer by rfer(16322)

38/6=57/x
38x=342
x=342/38
x=9
-----------
38-12=26
26/2=13
-----------------
57-18=39
39/2=19.5
------------
2*6+2*13=38
-------------
2*9+2*19.5=57